FANDOM


Apollo11EagleNotebookPg064
<Brief page summary to go here.>

Grumman Aircraft Engineering Corporation

LOG

Page No 73114

Engineer

Homburger / Hecht

Project

LM-5

Location

ACE 3

Time

Day Shift

Title

{{{Title}}}

Date

8-15-68

(1) Continued T/S on TDR #33. Had to power up for other subsystem testing after lunch. TDR #33 in hold.

(2) Notified by K. Leuchter that PSAPIM connector opened by EPS Seq 77 (X-lunar Bus Isol.) MUST BE CLOSED prior to Monday Morning for this test.

3. Had J722 to PTMU re-mated. Note -- OCP Seq. 077 call for P722 going into the EMI box to be de-mated for the test. Waiting for window to try this.

4. Powered-down via TDR #33 continuation sheet so that STE could then power-up via 70010 to give power to other systems.

5. Data taken yesterday on Ascent ECA's indicate high resistance insdie ECA #4 between terminals 4 and 2. (TDR Sheets 39 thru 42)

6. Try to complete TDR #33 as follows:

(a) Demate proper P722 conn & do not exceed 1 hr max time.
(b) If not successful, try demating vehicle side of Lut Interface Bob (at P92). (We heard rumors [sic: roumors] of sensing leads in Bob which are not shown on our schematic).
(c) If this is successful, complete the seq (077).

7. If time permits (write on TDR #3) remove all potting on ECA #4 (Bat 6) terminals #1, 2 & 4 in preparation for testing tomorrow. Note: latest TDR #3 & #33 sheets are in our black notebook.

8. Continued on TDR #33, page 4 at 6:20 pm after RCS completed. Stopped at 077-012.

9. Established that info on J722 remate (received fr. LoCassio) was wrong. Joe Palazzo, QC, had verified last night that P722 was [illegible -- bottom of page cut off].

Previous page (63) index Next page (65)